OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates (8, 3/2pi)

OpenStudy (science0229):

Do you know the relationship between the rectangular and polar coordinates?

OpenStudy (anonymous):

I know that you would use \[(r,\theta)=r(\cos \theta,\sin \theta)\]

OpenStudy (anonymous):

I got (0,-8) is that correct?

OpenStudy (science0229):

yeah!

OpenStudy (anonymous):

Thanks!